OpenStudy (anonymous):

Evaluate the follow integral and simplify the result. ∫1/((2e^x) -3)

OpenStudy (mimi_x3):

\[\int \frac{1}{2e^x-3}dx\]??

OpenStudy (anonymous):

Yes, can you help me?

OpenStudy (mimi_x3):

let \(u=e^x\) => du/dx=e^{x}

OpenStudy (mimi_x3):

\[\int \frac{1}{2u-3}*e^{x}du =>\int \frac{1}{u(2u-3)}du\]

OpenStudy (anonymous):

Wait, I thought that if du/dx=e^{x}, then dx=du/e^x.

OpenStudy (mimi_x3):

\[\frac{du}{dx}=e^x\] \[du=e^{x}dx\] \[e^{x}dx=du\] \[dx=\frac{du}{e^x}\]

OpenStudy (mimi_x3):

yeah, i made a typo..but the last part is still right.

OpenStudy (anonymous):

So is it \[\int\limits_{}^{}\frac{ 1 }{ 2e^x-3 } \frac{ du }{ e^x }\]

OpenStudy (mimi_x3):

u=e^x

OpenStudy (mimi_x3):

then proceed with partial fractions.